# Neural Conditional Transport Maps

### v1

@@ Appendix A Supplementary Material
  • In A.1, we provide extensive implementation details for our models, which should enhance reproducibility. Note that we will share our code upon acceptance.
@@ A.1 Implementation Details
Our implementation uses residual blocks for both $T_{\theta}$ and $f_{\omega}$. Each block consists of Layer Normalization [Ba et al., 2016], linear transformations with SiLU activation [Ramachandran et al., 2017], and residual connections with learnable scaling. Residual connections improve gradient flow in deep networks, while the learnable scaling parameter allows the network to control the contribution of the residual path during early training stages. Layer Normalization stabilizes training across a wide range of batch sizes and learning rates, particularly important in our adversarial setting.
We use orthogonal initialization [Saxe et al., 2014] and AdamW [Loshchilov and Hutter, 2019]. Orthogonal initialization preserves gradient norms through deep networks, addressing vanishing and exploding gradient problems typical in transport map training. Gradient norm clipping with threshold $1.0$ prevents gradient explosion during training, especially critical during the adversarial optimization process where critic and transport networks can destabilize each other.
In our experiments, following Korotin et al. [2023], we use the squared Euclidean cost: $k(T(x,z),y)=\|T(x,z)-y\|_{2}^{2}$, but other ground costs could be used. During pre-training, the loss $\mathcal{L}_{f}^{\text{pre}}$ uses equal weights ($\lambda_{\text{smooth}}=\lambda_{\text{transport}}=\lambda_{\text{mag}}=1.0$), with noise $\epsilon\sim\mathcal{N}(0,0.05)$ for the smoothness term. The smoothness term encourages local continuity in the critic, while the magnitude term prevents unbounded growth of critic values. Pre-training runs for 500 steps before transitioning to the alternating optimization of $T_{\theta}$ and $f_{\omega}$ for 5000 epochs, initializing the networks in favorable regions of the parameter space before adversarial training.
@@ A.2 Implementation Guidelines
Through extensive experimentation, we identified several consistent patterns for optimal neural conditional transport implementations. For network architectures, we observed that $T$ should have approximately $1.5-3$ times more trainable parameters than $f$ across all applications. This increased capacity should specifically be allocated through additional hidden layers rather than increasing layer width. Notably, $T$ benefits from an asymmetrical architecture with more decoder layers than encoder layers. This asymmetry aligns with the functional roles: the encoder performs the comparatively simpler task of feature extraction, while the decoder must simultaneously perform the transport mapping and appropriately integrate the conditioning information.
@@ A.3 Global Sensitivity Analysis
Global sensitivity analysis (GSA) studies how the uncertainty in the model output can be apportioned to different sources of uncertainty in the model input. Thus, GSA is crucial when developing and deploying complex models [Saltelli, 2002]. It enables users to understand the parametric components of the model, whether they are inputs like in machine learning models or parametric assumptions like in climate-economy models. OT has been recently introduced in the GSA literature by the works of Wiesel [2022] and Borgonovo et al. [2024]. Here, we present an overview of these OT-based sensitivity indices.
Let’s assume we want to represent some quantities of interest $\mathbf{Y}=(Y_{1},\dots,Y_{k})$ as a function of the inputs $\mathbf{C}=(C_{1},\dots,C_{d})$. We also assume that $\mathbf{C}$ and $\mathbf{Y}$ are random vectors on some probability space $(\Omega,\mathcal{B},\mu)$, and we define with $\mu_{C_{i}}$ and $\mu_{\mathbf{Y}}$ the distributions of $C_{i}$ and $\mathbf{Y}$, respectively. We consider a model defined by the function $\mathbf{f}\colon\mathcal{C}\subset\mathbb{R}^{d}\longrightarrow\mathbb{R}^{k}$. Using the OT cost in Equation (1), we can define the importance of the $i$-th input as:
@@
The rationale behind the index is simple. First, we fix the value of $C_{i}$ at $c_{i}$ and compute the conditional distribution of the output $\mathbf{Y}$ given this information, denoted as $\mu_{\mathbf{Y}|C_{i}=c_{i}}$. Second, we use the metric properties of the OT cost [Peyre and Cuturi, 2019, Chapter 2] to quantify the impact of fixing $C_{i}$ at $c_{i}$. As a third step, the index is computed as the expected value of the OT cost over the domain $\mathcal{C}$. Finally, everything is normalized by the upper bound $\mathbb{E}[k(\mathbf{Y},\mathbf{Y}^{\prime})]$.
The indices in Equation (7) have relevant properties such as zero-independence and max-functionality. Zero-independence reassures us that $\iota^{K}(\mathbf{Y},C_{i})\geq 0$ and $\iota^{K}(\mathbf{Y},C_{i})=0$ if and only if $\mathbf{Y}$ is independent to $C_{i}$, while Max-functionality entails that $\iota^{K}(\mathbf{Y},C_{i})\leq 1$ and $\iota^{K}(\mathbf{Y},C_{i})=1$ if and only if there exists a measurable function $\mathbf{g}$ such that $\mathbf{Y}=\mathbf{g}(C_{i})$.
@@
The first term in Equation (8) is the U-statistic of the upper bound, and we denote with $\mu^{N}$ the empirical distributions. In the original work, Borgonovo et al. [2024] suggest using well-established and fast solvers like network flow and transportation simplex [Luenberger and Ye, 2021] to compute $K(\mu^{N}_{\mathbf{Y}},\mu^{N}_{\mathbf{Y}|C_{i}\in\mathcal{C}_{i}^{m}})$. The first limitation is that their application requires the computation of the cost matrix, which can be highly memory-intensive for large $N$. Moreover, these solvers do not exploit the potential information in the partition ordering. In our approach, we compute the solution using Equation (5): $K(\mu^{N}_{\mathbf{Y}},\mu^{N}_{\mathbf{Y}|C_{i}\in\mathcal{C}_{i}^{m}})=\sup_{f}\inf_{T}\mathcal{L}(f,T,C_{i}\in\mathcal{C}_{i}^{m})$. Using the conditioned neural solver, we rely on the network structure to share information between partitions and avoid computing and storing the cost matrix.

### v2

@@ Appendix A Supplementary Material
  • In A.1, we provide extensive implementation details for our models, which should enhance reproducibility.
@@ A.1 Implementation Details
Our implementation uses residual blocks for both $T_{\theta}$ and $f_{\omega}$. Each block consists of Layer Normalization (Ba et al., 2016), linear transformations with SiLU activation (Ramachandran et al., 2017), and residual connections with learnable scaling. Residual connections improve gradient flow in deep networks, while the learnable scaling parameter allows the network to control the contribution of the residual path during early training stages. Layer Normalization stabilizes training across a wide range of batch sizes and learning rates, particularly important in our adversarial setting.
We use orthogonal initialization (Saxe et al., 2014) and AdamW (Loshchilov and Hutter, 2019). Orthogonal initialization preserves gradient norms through deep networks, addressing vanishing and exploding gradient problems typical in transport map training. Gradient norm clipping with threshold $1.0$ prevents gradient explosion during training, especially critical during the adversarial optimization process where critic and transport networks can destabilize each other.
In our experiments, following Korotin et al. (2023), we use the squared Euclidean cost: $k(T(x,z),y)=\|T(x,z)-y\|_{2}^{2}$, but other ground costs could be used. During pre-training, the loss $\mathcal{L}_{f}^{\text{pre}}$ uses equal weights ($\lambda_{\text{smooth}}=\lambda_{\text{transport}}=\lambda_{\text{mag}}=1.0$), with noise $\epsilon\sim\mathcal{N}(0,0.05)$ for the smoothness term. The smoothness term encourages local continuity in the critic, while the magnitude term prevents unbounded growth of critic values. Pre-training runs for 500 steps before transitioning to the alternating optimization of $T_{\theta}$ and $f_{\omega}$ for 5000 epochs, initializing the networks in favorable regions of the parameter space before adversarial training.
@@ A.2 Implementation Guidelines
Through extensive experimentation, we identified several consistent patterns for optimal neural conditional transport implementations. For network architectures, we observed that $T$ should have approximately 1.531.5-3 times more trainable parameters than $f$ across all applications. This increased capacity should specifically be allocated through additional hidden layers rather than increasing layer width. Notably, $T$ benefits from an asymmetrical architecture with more decoder layers than encoder layers. This asymmetry aligns with the functional roles: the encoder performs the comparatively simpler task of feature extraction, while the decoder must simultaneously perform the transport mapping and appropriately integrate the conditioning information.
@@ A.3 Global Sensitivity Analysis
Global sensitivity analysis (GSA) studies how the uncertainty in the model output can be apportioned to different sources of uncertainty in the model input. Thus, GSA is crucial when developing and deploying complex models (Saltelli, 2002). It enables users to understand the parametric components of the model, whether they are inputs like in machine learning models or parametric assumptions like in climate-economy models. OT has been recently introduced in the GSA literature by the works of Wiesel (2022) and Borgonovo et al. (2024). Here, we present an overview of these OT-based sensitivity indices.
Let’s assume we want to represent some quantities of interest $\mathbf{Y}=(Y_{1},\dots,Y_{k})$ as a function of the inputs $\mathbf{C}=(C_{1},\dots,C_{d})$. We also assume that $\mathbf{C}$ and $\mathbf{Y}$ are random vectors on some probability space $(\Omega,\mathcal{B},\mu)$, and we define with $\mu_{C_{i}}$ and $\mu_{\mathbf{Y}}$ the distributions of $C_{i}$ and $\mathbf{Y}$, respectively. We consider a model defined by the function $\mathbf{f}\colon\mathcal{C}\subset\mathbb{R}^{d}\longrightarrow\mathbb{R}^{k}$. Using the OT cost in equation 1, we can define the importance of the $i$-th input as:
@@
The rationale behind the index is simple. First, we fix the value of $C_{i}$ at $c_{i}$ and compute the conditional distribution of the output $\mathbf{Y}$ given this information, denoted as $\mu_{\mathbf{Y}|C_{i}=c_{i}}$. Second, we use the metric properties of the OT cost (Peyré and Cuturi, 2019, Chapter 2) to quantify the impact of fixing $C_{i}$ at $c_{i}$. As a third step, the index is computed as the expected value of the OT cost over the domain $\mathcal{C}$. Finally, everything is normalized by the upper bound $\mathbb{E}[k(\mathbf{Y},\mathbf{Y}^{\prime})]$.
The indices in equation 8 have relevant properties such as zero-independence and max-functionality. Zero-independence reassures us that $\iota^{K}(\mathbf{Y},C_{i})\geq 0$ and $\iota^{K}(\mathbf{Y},C_{i})=0$ if and only if $\mathbf{Y}$ is independent to $C_{i}$, while Max-functionality entails that $\iota^{K}(\mathbf{Y},C_{i})\leq 1$ and $\iota^{K}(\mathbf{Y},C_{i})=1$ if and only if there exists a measurable function $\mathbf{g}$ such that $\mathbf{Y}=\mathbf{g}(C_{i})$.
@@
The first term in equation 9 is the U-statistic of the upper bound, and we denote with $\mu^{N}$ the empirical distributions. In the original work, Borgonovo et al. (2024) suggest using well-established and fast solvers like network flow and transportation simplex (Luenberger and Ye, 2021) to compute $K(\mu^{N}_{\mathbf{Y}},\mu^{N}_{\mathbf{Y}|C_{i}\in\mathcal{C}_{i}^{m}})$. The first limitation is that their application requires the computation of the cost matrix, which can be highly memory-intensive for large $N$. Moreover, these solvers do not exploit the potential information in the partition ordering. In our approach, we compute the solution using Equation equation 5: $K(\mu^{N}_{\mathbf{Y}},\mu^{N}_{\mathbf{Y}|C_{i}\in\mathcal{C}_{i}^{m}})=\sup_{f}\inf_{T}\mathcal{L}(f,T,C_{i}\in\mathcal{C}_{i}^{m})$. Using the conditioned neural solver, we rely on the network structure to share information between partitions and avoid computing and storing the cost matrix.